OpenStudy (anonymous):

f(x)= (6x)−1/2 + 5 f ' (x) =?

OpenStudy (anonymous):

6x^-1/2

OpenStudy (anonymous):

(-1/2)(6)(x^(-3/2)

OpenStudy (lgbasallote):

-1/2(6x^-3/2) -1/[2 sqrt of (6x)^3]

OpenStudy (campbell_st):

well f'(x) = -3x^-3/2 \[f '(x) = -3/\sqrt{x^3}\]

OpenStudy (anonymous):

i've already tried -3x^-3/2 and -3/sqrt3 but i have not tried -1/2 sqrt of 6x^3

OpenStudy (lgbasallote):

nahhh. that was wrong :)) 6 should be in the numerator

OpenStudy (anonymous):

somehow the answer is -1/2x*sqrt6x..... dont know how

OpenStudy (campbell_st):

just look at my solution....

OpenStudy (anonymous):

idk how though

OpenStudy (campbell_st):

you have -1/2 * 6 = -3 x^(-3/2) = 1/x^3/2 so its 3 * 1/x^3/2